# Complexity of Finite Borel Asymptotic Dimension

### v1

@@ 3. Complexity of forward-independent hitting sets
In this section, we use a general condition for $\mathbf{\Sigma}^{1}_{2}$-completeness due to Frisch, Shinko, and Vidnyánszky [FSV24, Theorem 3.1] to deduce the following theorem, which immediately implies that the set (of codes) of Borel functions admitting Borel forward-independent hitting sets is $\mathbf{\Sigma}^{1}_{2}$-complete. Our main result, Theorem 1.1, is a corollary of Theorem 3.1 together with Theorem 1.2, which is proven in Section 5.
@@
with the property that, for every $\sigma\in\mathbb{N}^{\mathbb{N}}$, there is $f_{\sigma}:X\to X$ such that
@@
and such that
[…]
@@
[…]
@@
Figure 2. The structure $\mathcal{D}_{r}$ on $\mathbb{N}$ for various values of $r\in\mathbb{N}^{+}$.
@@
Now we are ready to prove Theorem 3.1. The proof strategy is as follows: We show first that, while there are no Borel forward-independent hitting sets for $S$ on the entirety of $[\mathbb{N}]^{\mathbb{N}}$, such sets can be constructed for $S$ on the non-dominating subsets of $[\mathbb{N}]^{\mathbb{N}}$, i.e., the sets of the form $\text{Dom}_{x}$ where $x\in[\mathbb{N}]^{\mathbb{N}}$; furthermore, this construction can be done in a uniform Borel way. This then allows us to apply [FSV24, Theorem 3.1], which gives the $\mathbf{\Sigma}^{1}_{2}$-completeness of the sets (3.2) and (3.3).
@@
We first fix $r\in\mathbb{N}^{+}$ and prove that there are a standard Borel space $X$ and a Borel function $f:X\to X$ as in the statement of Theorem 3.1 such that the set (3.2) is $\mathbf{\Sigma}^{1}_{2}$-complete. At the end of the proof, we explain how to re-select $X$ and $f$ so that the set (3.1) is $\mathbf{\Sigma}^{1}_{2}$-complete.
Recall that $\overrightarrow{\mathcal{G}}_{S}$ is the Borel digraph on $[\mathbb{N}]^{\mathbb{N}}$ induced by the shift function $S$ on $[\mathbb{N}]^{\mathbb{N}}$. Note that there is no Borel map $\varphi:[\mathbb{N}]^{\mathbb{N}}\to\mathbb{N}$ that is a homomorphism from $([\mathbb{N}]^{\mathbb{N}},\overrightarrow{\mathcal{G}}_{S})$ to $(\mathbb{N},\mathcal{D}_{r})$. This follows, for example, from [KST99, Example 3.2], since the existence of a Borel $r$-forward-independent hitting set for $S$ would imply that the Borel chromatic number of $\mathcal{G}_{S}$ is finite.
In order to apply [FSV24, Theorem 3.1], we must show that there is a Borel map $\varphi_{r}:\text{Dom}\to\mathbb{N}$ such that, for each $x\in[\mathbb{N}]^{\mathbb{N}}$, the restriction $\varphi_{r}\restriction\text{Dom}_{x}$ is a homomorphism from $\overrightarrow{\mathcal{G}}_{S}\restriction\text{Dom}_{x}$ to $\mathcal{D}_{r}$, where $\overrightarrow{\mathcal{G}}_{S}\restriction\text{Dom}_{x}$ denotes the sub-digraph of $\overrightarrow{\mathcal{G}}_{S}$ induced by $\text{Dom}_{x}$.
Set $x(-r)=0$ for all $x\in[\mathbb{N}]^{\mathbb{N}}$. Define $\mathcal{H}_{r}\subseteq([\mathbb{N}]^{\mathbb{N}})^{2}$ by $(x,y)\in\mathcal{H}_{r}$ if and only if $|[x(2rn-r),x(2rn+r))\cap y|\equiv_{2r}0$, where $n\in\mathbb{N}$ is minimal such that $[x(2rn-r),x(2rn+r))\cap y\neq\emptyset$. Observe that $\mathcal{H}_{r}$ is Borel.
@@
If $(x,y)\in\text{Dom}$, then there is $n\in\mathbb{N}$ such that $|[x(2rn-r),x(2rn+r)\cap y|\geq 2r$.
@@
where $n\in\mathbb{N}$ is minimal such that $[x(2rn-r),x(2rn+r))\cap y\neq\emptyset$. If $\varphi_{r}(x,y)>0$, then $\varphi_{r}(x,S(y))=\varphi_{r}(x,y)-1$ by the definition.
Now we are ready to finish this part of the proof. Note that the assumptions of [FSV24, Theorem 3.1] are satisfied by the construction of $\varphi_{r}$ above. Hence, there is a Borel set $B\subseteq\mathbb{N}^{\mathbb{N}}\times\mathbb{N}^{\mathbb{N}}\times[\mathbb{N}]^{\mathbb{N}}$ such that
(3.3) $\{\sigma\in\mathbb{N}^{\mathbb{N}}:\text{ there is a Borel homomorphism from }(\overrightarrow{\mathcal{G}}^{\prime}_{S}\restriction B_{\sigma})\text{ to }\mathcal{D}_{r}\}$
@@
and observe that $Y$ is a Borel set. Define $X^{\prime}=B\sqcup\bigsqcup_{\ell\in\mathbb{N}^{+}}Y^{\ell}$, where $Y^{\ell}=Y\times\{\ell\}$, and set $f(\sigma,\gamma,x)=(\sigma,\gamma,S(x))$ whenever $(\sigma,\gamma,x)\in B\setminus Y$, $f(y)=(y,1)$ for $y\in Y$ and $f((y,\ell)))=(y,\ell+1)$ for $(y,\ell)\in Y^{\ell}$. It follows that $f:X^{\prime}\to X^{\prime}$ is an acyclic Borel function and, using the Borel isomorphism theorem, we may assume that $X^{\prime}$ is of the form ×$\mathbb{N}^{\mathbb{N}}\times X$ for some standard Borel space $X$. It is easy to see, using for example Remark 1.3, that the sets (3.3) and (3.2) coincide.
We now explain how to modify the proof to obtain $X$ and $f$ so that (3.1) is $\mathbf{\Sigma}^{1}_{2}$-complete: Consider a language consisting of countably many binary relation symbols $(\mathcal{R}_{r})_{r\in\mathbb{N}^{+}}$. For each $r\in\mathbb{N}^{+}$, interpret $\mathcal{R}_{r}$ in $\mathbb{N}$ as the edge relation of $\mathcal{D}_{r}$, and apply [FSV24, Theorem 3.1] to obtain a Borel set $B\subseteq\mathbb{N}^{\mathbb{N}}\times\mathbb{N}^{\mathbb{N}}\times[\mathbb{N}]^{\mathbb{N}}$ such that
(3.4) $\{\sigma\in\mathbb{N}^{\mathbb{N}}:\forall r\in\mathbb{N}^{+}\text{ there is aBorel homomorphism from }(\overrightarrow{\mathcal{G}}^{\prime}_{S}\restriction B_{\sigma})\text{ to }\mathcal{D}_{r}\}$
@@
In personal communication, Alex Kastner pointed out that the Borel set $B\subseteq\mathbb{N}^{\mathbb{N}}\times\mathbb{N}^{\mathbb{N}}\times[\mathbb{N}]^{\mathbb{N}}$ obtained from an application of [FSV24, Theorem 3.1] can be chosen so that, for each $\sigma,\gamma\in\mathbb{N}^{\mathbb{N}}$, the section $B_{\sigma,\gamma}$ is $\mathcal{G}_{S}$-invariant. This allows one to define $X^{\prime}=B$ and $f(\sigma,\gamma,x)=(\sigma,\gamma,S(x))$ directly in the proof of Theorem 3.1.
@@ 4. The classification of $\operatorname{CSP}^{\operatorname{function}}_{B}(H)$
This section is devoted to a proof of Theorem 1.4. As in the proof of of Theorem 3.1, we utilize [FSV24, Theorem 3.1].
Recall that the set of codes for Borel functions on a fixed standard Borel space $X$ is $\mathbf{\Pi}^{1}_{1}$ [HKM24, Lemma A.3]. For the remainder of the section, we fix a finite sinkless digraph $H$.
@@
(ii). Assume that $H$ is an ergodic digraph with no loops. Then by the previous paragraph, we have $\overrightarrow{\mathcal{G}}_{S}\not\in\operatorname{CSP}^{\operatorname{function}}_{B}(H)$. The next result, Proposition 4.1, guarantees that [FSV24, Theorem 3.1] can be used as in the proof of Theorem 3.1 to obtain an acyclic Borel function $f=(f_{\sigma})_{\sigma\in{\mathbb{N}^{\mathbb{N}}}}:\mathbb{N}^{\mathbb{N}}\times X\to\mathbb{N}^{\mathbb{N}}\times X$ such that
@@
By Claim 4.2, we may, without loss of generality, assume that $H$ is a disjoint union of strongly connected digraphs. In particular, $H$ is both sinkless and sourceless. By the assumption in (iii), none of the strong connectivity components of $H$ is ergodic; then by [Tho22, Corollary 6.3], $\operatorname{CSP}_{B}(H)$ is $\mathbf{\Pi}^{1}_{1}$. Consequently, $\operatorname{CSP}^{\operatorname{function}}_{B}(H)$ is $\mathbf{\Pi}^{1}_{1}$ as well. ∎
@@ 5. Finite asymptotic dimension for graphs generated by a Borel function
In this section, we prove Theorems 1.2 and 1.1. We begin by recalling the definition of Borel asymptotic dimension for a Borel graph [CJM+23, Definition 3.2]. In [CJM+23], two separate definitions – one in terms of coverings and the other in terms of equivalence relations – are provided; these are shown to be equivalent for locally countable Borel graphs [CJM+23, Lemma 3.1]. We prove that, for graphs generated by a single Borel function $f$, the two definitions remain equivalent even when $f$ is uncountable-to-one.
Fix a standard Borel space $X$. Since we work exclusively with graphs of the form $\mathcal{G}_{f}$, where $f$ is a Borel function on $X$, we denote by $\rho_{f}$ the graph distance metric of $\mathcal{G}_{f}$. Given $t\in\mathbb{N}^{+}$ and $x\in X$, we write $B_{t}(x)$ for the $\rho_{f}$-ball of radius $t$ around $x\in X$. Given $U\subseteq X$ and $r\in\mathbb{N}^{+}$, we define $\mathcal{F}_{r}(U)$ to be the equivalence relation on $U$ generated by the set of pairs $(x,y)\in U^{2}$ such that $\rho_{f}(x,y)\leq r$.
@@ 5.1. From finite asymptotic dimension to hitting sets
Then $A$ is Borel hitting set for $f$. Indeed, we have that $A=\bigcap_{k\in\mathbb{N}^{+}}A_{k}$, where
@@
is Borel for each $k\in\mathbb{N}^{+}$ by the Lusin–Suslin theorem. Let $x\in X$. Then since $[x]_{E}$ has finite diameter, there is $k\in\mathbb{N}$ such that $(x,f^{k}(x))\in E$ and $(x,f^{k+j}(x))\notin E$ for every $j\in\mathbb{N}^{+}$. Thus $f^{k}(x)\in A$.
@@
Then $H=A\setminus(\bigcup_{1\leq j\leq t}f^{-j}(A))$ is Borel and $t$-forward-independent. We claim that $H$ is hitting. Let $x_{0}\in X$. Since $A$ is hitting, there is $x\in H$ that is an $f$-iterate of $x_{0}$. If $x\in H$, then there is nothing to prove. Otherwise, there is $1\leq j_{0}\leq r_{0}$ such that $f^{j_{0}}(x)\in A$. Repeat this procedure $d+1$ times to obtain $1\leq j_{0},\dots,j_{d}\leq r_{0}$ such that either $f^{j_{0}+\cdots+j_{m}}(x)\in H$ for some $m\leq d$ or $f^{j_{0}+\cdots+j_{m}}(x)\in A$ for each $m\leq d$. Note that $B_{2t(d+1)}(x)$ contains $f^{j_{0}+\cdots+j_{m}}(x)\in A$ for each $m\leq d$. If $m\neq m^{\prime}$, then
$(f^{j_{0}+\cdots+j_{m}}(x),f^{j_{0}+\cdots+j_{m^{\prime}}}(x))\notin E,$
and $(x,f^{j_{0}+\cdots+j_{m}}(x))\notin E$ for any $m\leq d$. But then $B_{r}(x)$ meets at least $(d+2)$ $E$-classes, a contradiction. ∎
@@
Note that a uniform bound on diameter is not needed to construct the hitting sets in these proofs; boundedness alone is sufficient. Therefore, one can show that, if $\mathcal{G}_{f}$ has finite Borel asymptotic separation index, denoted $\text{asi}_{B}(\mathcal{G}_{f})$ (see [CJM+23, Definition 3.2]), then there is a Borel $r$-forward-independent hitting set for $f$ for each $r\in\mathbb{N}^{+}$. In particular, for a graph of the form $\mathcal{G}_{f}$, if $\text{asi}_{B}(\mathcal{G}_{f})<\infty$, then $\text{asi}_{B}(\mathcal{G}_{f})\leq 1$.

### v2

@@ 3. Complexity of forward-independent hitting sets
In this section, we use Theorem 2.3 to deduce the following theorem, which immediately implies that the set (of codes) of Borel functions admitting Borel forward-independent hitting sets is $\mathbf{\Sigma}^{1}_{2}$-complete. Our main result, Theorem 1.1, is a corollary of Theorem 3.1 together with Theorem 1.2, which is proved in Section 5.
@@
with the property that, for every $\sigma\in\mathbb{N}^{\mathbb{N}}$, there exists $f_{\sigma}:X\to X$ such that
@@
and
(3.2) $\{\sigma\in\mathbb{N}^{\mathbb{N}}:\forall r\in\mathbb{N}^{+}\ f_{\sigma}\ \text{admits a Borel }r\text{-forward-independent hitting set}\}$
@@
(3.3) $\{\sigma\in\mathbb{N}^{\mathbb{N}}:f_{\sigma}\ \text{admits a Borel }r\text{-forward-independent hitting set}\}$
@@
Figure 2. The structure $\mathcal{D}_{r}$ on $\mathbb{N}$ for various values of $r\in\mathbb{N}^{+}$
@@
Now we are ready to prove Theorem 3.1. The proof strategy is as follows: We show first that, while there are no Borel forward-independent hitting sets for $S$ on the entirety of $\mathbb{N}$$^{\mathbb{N}}$, such sets can be constructed for $S$ on the non-dominating subsets of $\mathbb{N}$$^{\mathbb{N}}$, i.e., the sets of the form $\text{Dom}_{x}$ where $x\in[\mathbb{N}]^{\mathbb{N}}$; furthermore, this construction can be done in a uniform Borel way. This then allows us to apply [18, Theorem 3.1], which gives the $\mathbf{\Sigma}^{1}_{2}$-completeness of the sets (3.2) and (3.3).
@@
We first fix $r\in\mathbb{N}^{+}$ and prove that there are a standard Borel space $X$ and a Borel function $f:\mathbb{N}^{\mathbb{N}}\times X\to\mathbb{N}^{\mathbb{N}}\times X$ as in the statement of Theorem 3.1 such that the set (3.3) is $\mathbf{\Sigma}^{1}_{2}$-complete. At the end of the proof, we explain how to re-select $X$ and $f$ so that the set (3.2) is $\mathbf{\Sigma}^{1}_{2}$-complete.
Recall that $\overrightarrow{\mathcal{G}}_{S}$ is the Borel digraph on $\mathbb{N}$$^{\mathbb{N}}$ induced by the shift function $S$ on $\mathbb{N}$$^{\mathbb{N}}$. Note that there is no Borel map $\varphi:[\mathbb{N}]^{\mathbb{N}}\to\mathbb{N}$ that is a homomorphism from $([\mathbb{N}]^{\mathbb{N}},\overrightarrow{\mathcal{G}}_{S})$ to $(\mathbb{N},\mathcal{D}_{r})$. This follows, for example, from [28, Example 3.2], since the existence of a Borel $r$-forward-independent hitting set for $S$ would imply that the Borel chromatic number of $\mathcal{G}_{S}$ is finite.
In order to apply Theorem 2.3, we must show that there is a Borel map $\varphi_{r}:\text{Dom}\to\mathbb{N}$ such that, for each $x\in[\mathbb{N}]^{\mathbb{N}}$, the restriction $\varphi_{r}\restriction\text{Dom}_{x}$ is a homomorphism from $\overrightarrow{\mathcal{G}}_{S}\restriction\text{Dom}_{x}$ to $\mathcal{D}_{r}$, where $\overrightarrow{\mathcal{G}}_{S}\restriction\text{Dom}_{x}$ denotes the sub-digraph of $\overrightarrow{\mathcal{G}}_{S}$ induced by $\text{Dom}_{x}$.
Set x(r)=0x(-r)=0 for all $x\in[\mathbb{N}]^{\mathbb{N}}$. Define $\mathcal{H}_{r}\subseteq([\mathbb{N}]^{\mathbb{N}})^{2}$ by $(x,y)\in\mathcal{H}_{r}$ if and only if $|[x(2rn-r),x(2rn+r))\cap y|\equiv_{2r}0$, where $n\in\mathbb{N}$ is minimal such that $x(2rn-r),x(2rn+r))\cap y\neq\emptyset$. Observe that $\mathcal{H}_{r}$ is Borel.
@@
If $(x,y)\in\text{Dom}$, then there is $n\in\mathbb{N}$ such that $|[x(2rn-r),x(2rn+r))\cap y|\geq 2r$.
@@
where $n\in\mathbb{N}$ is minimal such that $x(2rn-r),x(2rn+r))\cap y\neq\emptyset$. If $\varphi_{r}(x,y)>0$, then $\varphi_{r}(x,S(y))=\varphi_{r}(x,y)-1$ by the definition.
Now we are ready to finish the first part of the proof. Note that the assumptions of Theorem 2.3 are satisfied by the construction of $\varphi_{r}$ above. Hence, there is a Borel set $B\subseteq\mathbb{N}^{\mathbb{N}}\times\mathbb{N}^{\mathbb{N}}\times[\mathbb{N}]^{\mathbb{N}}$ such that
[…]
@@
and observe that $Y$ is a Borel set. Set also $Z=(\mathbb{N}^{\mathbb{N}}\times\mathbb{N}^{\mathbb{N}}\times[\mathbb{N}]^{\mathbb{N}})\setminus B$. We define
[…]
and a function $f$ on $\mathbb{N}^{\mathbb{N}}\times X$ as follows. If $(\sigma,(\gamma,x,\ell))\in\mathbb{N}^{\mathbb{N}}\times X$ has the property that $(\sigma,\gamma,x)\not\in Y\cup Z$, then we set $f(\sigma,(\gamma,x,\ell))=(\sigma,(\gamma,S(x),\ell))$. Otherwise, we have that $(\sigma,\gamma,x)\in Y\cup Z$ and we let $f(\sigma,(\gamma,x,\ell))=(\sigma,(\gamma,x,\ell+1))$. Then $f:\mathbb{N}^{\mathbb{N}}\times X\to\mathbb{N}^{\mathbb{N}}\times X$ is a finite-to-one acyclic Borel function. Fix $\sigma\in\mathbb{N}^{\mathbb{N}}$, and let $f_{\sigma}$ be the function on $X$ such that $f(\sigma,(\gamma,x,\ell))=(\sigma,f_{\sigma}(\gamma,x,\ell))$ for all $(\gamma,x,\ell)\in X$. Note that the collection $\{f_{\sigma}\}_{\sigma\in\mathbb{N}^{\mathbb{N}}}$ is well-defined.
To complete this part of the proof, we show that the sets (3.6) and (3.3) coincide. Define a set $T\subseteq X$ by $(\gamma,x,\ell)\in T$ if $\ell=0$ and $(\sigma,\gamma,x)\in Y\cup Z$, and let $\mathcal{C}$ denote the collection of connected components of the graph $\mathcal{G}_{f_{\sigma}}$ on $X$ generated by $f_{\sigma}$ containing a point $(\gamma,x,\ell)\in X$ such that $(\sigma,\gamma,x)\in Y\cup Z$. Then $T$ is a Borel transversal for the connectedness relation of $\mathcal{G}_{f_{\sigma}}\restriction\bigcup\mathcal{C}$. Therefore, since $f_{\sigma}$ is acyclic, we have that $f_{\sigma}\restriction\bigcup\mathcal{C}$ admits a Borel $r$-forward-independent hitting set by Remark 1.3. It follows that the sets (3.6) and (3.3) coincide, as required.
We now explain how to modify the proof to obtain $X$ and $f$ so that (3.2) is $\mathbf{\Sigma}^{1}_{2}$-complete: Consider a language consisting of countably many binary relation symbols $(\mathcal{R}_{r})_{r\in\mathbb{N}^{+}}$. For each $r\in\mathbb{N}^{+}$, interpret $\mathcal{R}_{r}$ in $\mathbb{N}$ as the edge relation of $\mathcal{D}_{r}$, and apply Theorem 2.3 to obtain a Borel set $B\subseteq\mathbb{N}^{\mathbb{N}}\times\mathbb{N}^{\mathbb{N}}\times[\mathbb{N}]^{\mathbb{N}}$ such that
[…]
@@
In personal communication, Alex Kastner pointed out that the Borel set $B\subseteq\mathbb{N}^{\mathbb{N}}\times\mathbb{N}^{\mathbb{N}}\times[\mathbb{N}]^{\mathbb{N}}$ obtained from the application of Theorem 2.3 can be chosen so that, for each $\sigma,\gamma\in\mathbb{N}^{\mathbb{N}}$, the section $B_{\sigma,\gamma}$ is $\mathcal{G}_{S}$-invariant. This allows one to define $X=\mathbb{N}^{\mathbb{N}}\times[\mathbb{N}]^{\mathbb{N}}$ and $f(\sigma,\gamma,x)=(\sigma,\gamma,S(x))$ directly in the proof of Theorem 3.1.
@@ 4. The classification of $\operatorname{CSP}^{\operatorname{function}}_{B}(H)$
This section is devoted to a proof of Theorem 1.4. As in the proof of of Theorem 3.1, we utilize [18, Theorem 3.1].
Recall that the set of codes for Borel functions on a fixed standard Borel space $X$ is $\mathbf{\Pi}^{1}_{1}$ [23, Lemma A.3]. For the remainder of the section, we fix a finite sinkless digraph $H$.
@@
(ii). Assume that $H$ is an ergodic digraph with no loops. Then by the previous paragraph, we have $\overrightarrow{\mathcal{G}}_{S}\not\in\operatorname{CSP}^{\operatorname{function}}_{B}(H)$. The next result, Proposition 4.2, guarantees that [18, Theorem 3.1] can be used as in the proof of Theorem 3.1 to obtain an acyclic Borel function $f=(f_{\sigma})_{\sigma\in{\mathbb{N}^{\mathbb{N}}}}:\mathbb{N}^{\mathbb{N}}\times X\to\mathbb{N}^{\mathbb{N}}\times X$ such that
@@
By Claim 4.3, we may, without loss of generality, assume that $H$ is a disjoint union of strongly connected digraphs. In particular, $H$ is both sinkless and sourceless. By the assumption in (iii), none of the strong connectivity components of $H$ is ergodic; then by [40, Corollary 6.3], $\operatorname{CSP}_{B}(H)$ is $\mathbf{\Pi}^{1}_{1}$. Consequently, $\operatorname{CSP}^{\operatorname{function}}_{B}(H)$ is $\mathbf{\Pi}^{1}_{1}$ as well. ∎
@@ 5. Finite asymptotic dimension for graphs generated by a Borel function
In this section, we prove Theorems 1.2 and 1.1. We begin by recalling the definition of Borel asymptotic dimension for a Borel graph [10, Definition 3.2]. In [10], two separate definitions – one in terms of coverings and the other in terms of equivalence relations – are provided; these are shown to be equivalent for locally countable Borel graphs [10, Lemma 3.1]. We prove that, for graphs generated by a single Borel function $f$, the two definitions remain equivalent even when $f$ is uncountable-to-one.
Fix a standard Borel space $X$. Since we work exclusively with graphs of the form $\mathcal{G}_{f}$, where $f$ is a Borel function on $X$, we denote by $\rho_{f}$ the graph distance metric of $\mathcal{G}_{f}$. Given $t\in\mathbb{N}^{+}$ and $x\in X$, we write $B_{t}(x)$ for the $\rho_{f}$-ball of radius $t$ around $x\in X$. Given $U\subseteq X$ and $r\in\mathbb{N}^{+}$, we define $\mathcal{F}_{r}(U)$ to be the equivalence relation on $U$ generated by the set of pairs $(x,y)\in U^{2}$ such that $\rho_{f}(x,y)\leq r$. Given a Borel equivalence relation $E$ on $X$, we say $E$ is uniformly $\rho_{f}$-bounded (or merely uniformly bounded if $\rho_{f}$ is understood) if there is $M>0$ such that each $E$-class has $\rho_{f}$-diameter at most $M$.
@@ 5.1. From finite asymptotic dimension to hitting sets
Then $A$ is a Borel hitting set for $f$. Indeed, we have that $A=\bigcap_{k\in\mathbb{N}^{+}}A_{k}$, where
@@
is Borel for each $k\in\mathbb{N}^{+}$ by the Lusin–Suslin theorem. Now let $x\in X$. Then since $x]_{E}$ has finite diameter, there is $k\in\mathbb{N}$ such that $(x,f^{k}(x))\in E$ and $(x,f^{k+j}(x))\notin E$ for every $j\in\mathbb{N}^{+}$. Thus $f^{k}(x)\in A$.
@@
Then $H=A\setminus(\bigcup_{1\leq j\leq t}f^{-j}(A))$ is Borel and $t$-forward-independent. We claim that $H$ is hitting. Assume for contradiction that there is $x_{0}\in X$ such that $f^{k}(x_{0})\notin H$ for any $k\in\mathbb{N}$. Since $A$ is hitting, there are $x\in A$ and $\ell\in\mathbb{N}^{+}$ such that $f^{\ell}(x_{0})=x$. However, since $x\notin H$, there is $1\leq j_{1}\leq t$ such that $f^{j}(x)\in A$. Let $j_{0}=0$, and then recursively construct a sequence $(j_{n})_{n\geq 2}$ such that, for all $n\in\mathbb{N}$, $1\leq j_{n}\leq t$ and $f^{j_{1}+\cdots+j_{n}}(x)\in A$. Note that, for any $m\neq m^{\prime}$,
$(f^{j_{0}+\cdots+j_{m}}(x),f^{j_{0}+\cdots+j_{m^{\prime}}}(x))\notin E.$
So $B_{2t(d+1)}(x)$ meets at least $d+2$ classes of $E$, a contradiction. ∎
@@
Note that a uniform bound on diameter is not needed to construct the hitting sets in these proofs; boundedness alone is sufficient. Therefore, one can show that, if $\mathcal{G}_{f}$ has finite Borel asymptotic separation index, denoted $\text{asi}_{B}(\mathcal{G}_{f})$ (see [10, Definition 3.2]), then there is a Borel $r$-forward-independent hitting set for $f$ for each $r\in\mathbb{N}^{+}$. In particular, for a graph of the form $\mathcal{G}_{f}$, if $\text{asi}_{B}(\mathcal{G}_{f})<\infty$, then $\text{asi}_{B}(\mathcal{G}_{f})\leq 1$. It is open whether $\text{asi}_{B}(\mathcal{G})<\infty$ implies $\text{asi}_{B}(\mathcal{G})\leq 1$ for general Borel graphs $\mathcal{G}$ (see [10, 39, 2]).